### v1

# CamReasoner: Reinforcing Camera Movement Understanding via Structured Spatial Reasoning

Abstract
/Understanding camera dynamics is a fundamental pillar of video spatial intelligence. However, existing multimodal models predominantly treat this task as a black-box classification, often confusing physically distinct motions by relying on superficial visual patterns rather than geometric cues. We present CamReasoner, a framework that reformulates camera movement understanding as a structured inference process to bridge the gap between perception and cinematic logic. Our approach centers on the Observation-Thinking-Answer (O-T-A) paradigm, which compels the model to decode spatio-temporal cues such as trajectories and view frustums within an explicit reasoning block. To instill this capability, we construct a Large-scale Inference Trajectory Suite comprising 18k SFT reasoning chains and 38k RL feedback samples. Notably, we are the first to employ RL for logical alignment in this domain, ensuring motion inferences are grounded in physical geometry rather than contextual guesswork. By applying Reinforcement Learning to the Observation-Think-Answer (O-T-A) reasoning paradigm, CamReasoner effectively suppresses hallucinations and achieves state-of-the-art performance across multiple benchmarks.
[…]
1 Introduction
Camera movement serves as the invisible brushstroke of cinematic storytelling, acting as a vital conduit to convey subtle emotions and orchestrate audience immersion. Whether through Scorsese’s seamless tracking shots that create a seductive flow, or Spielberg’s dynamic dollies that heighten suspense and awe, these trajectories represent the intentionality of a director shaping the viewer’s perception of the physical and emotional world. In the landscape of Multimodal Large Language Models (MLLMs), camera movement understanding has emerged as a critical frontier that bridges the gap between passive pixel perception and active semantic reasoning. From the perspective of understanding, mastering these dynamics allows MLLMs to effectively decouple complex object movements from the observer’s ego-motion, even when camera intrinsic parameters are unknown, thereby enabling a more robust comprehension of 3D spatial geometry and narrative intent. This transition from surface-level recognition to deep structural analysis transforms an AI from a mere observer into a cinematic reasoner capable of interpreting the "why" behind every frame. Furthermore, from the perspective of generation, camera movement acts as the essential cornerstone for controllable video synthesis. As the frontier of AI shifts toward purposeful visual artistry, the ability to generate physically consistent videos relies fundamentally on the model’s grasp of how trajectories interact with scene depth and perspective. By integrating these motion priors, MLLMs can move beyond disorganized pixel synthesis toward achieving true world modeling capabilities, where every shift in perspective is both physically grounded and narratively coherent.
Despite the growing demand for such sophisticated motion perception, current methodologies for camera movement understanding remain split between two constrained paradigms that struggle to meet these requirements. Traditional geometric approaches, such as SfM [schonberger2016structure, wang2024vggsfm, zhang2023protein] and SLAM [davison2007monoslam, Engel2014LSD], rely on per-frame camera pose estimation to reconstruct trajectories, yet often require prior knowledge of camera intrinsics to maintain accuracy. These systems are fragile in dynamic environments because they often fail to disentangle the camera’s ego-motion from the complex movements of subjects within the scene. In contrast, modern learning-based Video-Language Models (VLMs) [comanici2025gemini, qwen3vl, hurst2024gpt] offer a promising alternative by treating motion understanding as a semantic task. Leveraging vast pre-training corpora, these models achieve impressive alignment between visual cues and cinematic terminology, while even demonstrating an emerging capacity for high-level reasoning [liu2025shotbench]. However, these models predominantly operate as black boxes that leap from pixels to labels without explicit reasoning. This shortcut often triggers hallucinations because the model relies on contextual cues like a running athlete instead of decoding the actual physics of the lens. Consequently, they struggle to distinguish physically distinct yet visually similar operations such as a zoom versus a dolly. These limitations necessitate a framework that bridges the gap between raw perception and structured cinematic logic.
Figure 1: Overview of CamReasoner. We propose the <observation>-<think>-<answer>paradigm to improve camera movement understanding. The figure illustrates our model generating detailed visual observations and logical thinking for movements like truck left and pan right.
To bridge this gap, we introduce CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task. Central to our approach is the Observation-Thinking-Answer (O-T-A) paradigm, which we implement through a sequential training pipeline of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). To support this process, we constructed a Large-scale Inference Trajectory Suite comprising 18k SFT and 38k RL samples. While the SFT data transforms static labels into explainable reasoning chains to instill cinematic logic, the RL samples enforce logical alignment to ground final verdicts in 3D spatio-temporal physics. By explicitly tracing trajectories and analyzing view frustums, our model establishes a rigorous logical premise before delivering a verdict, effectively eliminating contextual guesswork. As the first to apply Reinforcement Learning for logical alignment in this domain, CamReasoner-7B achieves state-of-the-art results, reaching 78.4% in binary classification and 74.5% in VQA. Notably, it excels in challenging Confusable Motion scenarios (60.7%) by distinguishing physical movements like truck versus pan through geometric reasoning. By integrating structural motion priors, CamReasoner provides a scalable path for MLLMs to achieve precise cinematic control and physically-consistent world modeling.
Our main contributions are summarized as follows:
  • Structured Reasoning Paradigm: We propose the Observation-Thinking-Answer (O-T-A) paradigm, reformulating camera movement understanding from black-box classification into a structured ego-motion inference task that decouples camera trajectories from dynamic scenes.
  • Inference Trajectory Suite: We construct a Large-scale Inference Trajectory Suite with 18k SFT and 38k RL samples, transforming static labels into dense reasoning chains to instill spatio-temporal cinematic logic into MLLMs.
  • RL-driven Logical Alignment: We are the first to employ Reinforcement Learning for logical alignment in this domain, achieving state-of-the-art 78.4% and 74.5% accuracy in binary classification and visual question answering tasks.
2 Related Work
2.1 Camera Movement Understanding
Camera movement understanding bridges geometric perception and semantic reasoning, yet current methods face significant bottlenecks. Traditional geometric approaches rely on SfM-based pseudo ground-truth, which is largely restricted to static scenes and ignores scene context [zhou2018stereo, ling2024dl3dv, jin2024stereo4d, schonberger2016structure]. Categorical recognition tasks, while providing semantic labels, often suffer from poor specification—confusing rotation with translation (e.g., pan vs. truck) and failing to capture co-occurring motions in complex cinematography [huang2020movienet, rao2020shot, argaw2022anatomy]. Although recent benchmarks like ShotBench [liu2025shotbench, wu2025refineshot] evaluate nuanced cinematic grammar, existing methodologies lack a structured reasoning process to explain the visual cues driving their interpretations [hong2025motion, wang2024tarsier]. Consequently, there remains a critical need for a framework that integrates spatial evidence with logical deduction to achieve a holistic understanding across diverse video domains [zhou2018stereo, huang2020movienet, argaw2022anatomy, liu2025shotbench].
2.2 Multimodal Reinforcement Learning
Inspired by the reasoning success of LLMs, recent works have extended structured thinking to MLLMs across diverse visual tasks [li2025star, sun2025reinforcement, feng2025video, sun2025spacevista, zhou2025reinforced, duan2025codeplot, chen2025advancing, chen2025ares, meng2025open]. These include task-specific reasoning models for image VQA [huang2025vision, mei2025surveycontextengineeringlarge, mei2025a1], video understanding [feng2025video, li2025videochat], object detection [yu2025perception, shen2025vlm], segmentation [you2025seg], and temporal grounding [wang2025time]. Notably, frameworks like SophiaVL-R1 [fan2025sophiavl], OneThinker [feng2025onethinker], and FrameMind [ge2025framemind] introduce unified reasoning via reinforcement learning, while Thinking-with-Sound [xiong2025thinking] extends this paradigm to the audio domain through acoustic tool manipulation. However, despite the surge in RL-based multimodal reasoning, its application to the specialized domain of camera movement understanding remains unexplored.
3 Method
We introduce a two-stage training strategy combining an SFT cold start with Group Relative Policy Optimization (GRPO) [shao2024deepseek] to instill the O-T-A reasoning paradigm into camera movement understanding. This pipeline ensures the model efficiently acquires structured reasoning traces while maintaining format adherence and task accuracy. To support this, we curate distinct datasets specifically tailored for sequential Supervised Fine-Tuning and Reinforcement Learning phases.
3.1 Data Collection and Curation
High-quality, structured data is pivotal for teaching the model cinematic logic. We constructed two specialized datasets to facilitate the O-T-A paradigm across the SFT and RL stages.
3.1.1 CamReasoning-SFT-18k Construction.
To instill structured reasoning, we curate the CamReasoning-SFT-18k dataset via an answer-conditioned synthesis pipeline. As shown in Fig. 2, we utilize Qwen2.5-VL-72B to generate 38,672 initial trajectories from original video-QA pairs, enforcing the <observation>, <think>, and <answer> format. To ensure logical coherence, we implement a multi-dimensional filtration process using Qwen3 [qwen3vl] to verify samples based on format adherence, label accuracy, and reasoning consistency. This results in 18,541 high-fidelity samples that serve as cold-start data for teaching the model step-by-step cinematic inference.
3.1.2 CamReasoning-RL-38k
For the reinforcement learning stage, we construct the CamReasoning-RL-38k dataset to drive GRPO [shao2024deepseek]. Derived from the CameraBench training split, this dataset provides a diverse array of camera movement scenarios converted into our standardized instruction schema. Unlike the SFT phase, these samples lack reasoning annotations; instead, they serve as environmental queries for the model to refine its self-generated reasoning paths. This setup allows the model to optimize logical robustness and prediction accuracy through trial and error, guided by task-specific reward signals.
Figure 2: Data generation pipeline for CamReasoning-SFT-18k. We utilize an answer-conditioned process to generate 38,672 initial trajectories, which are then filtered for format, accuracy, and consistency to retain 18,541 high-quality samples.
3.2 Training Strategy
Our training pipeline comprises two progressive stages designed to endow the model with robust O-T-A reasoning capabilities for camera movement understanding. We first employ SFT as a cold-start initialization to adapt the model to the structured <observation>-<think>-<answer>format. Subsequently, we advance to a RL phase utilizing GRPO to further refine the reasoning logic and align outputs with task requirements. Crucially, to mitigate the optimization instability inherent in single-task RL, we incorporate the EMA-GRPO strategy for robust advantage normalization.
3.2.1 Supervised Fine-tuning
We initiate the training process with SFT to bootstrap the model’s capability in generating structured O-T-A (Observation-Think-Answer) reasoning traces. As pre-trained MLLMs lack exposure to our specialized reasoning format encompassing <observation>, <think>, and <answer> tags, SFT provides the essential cold-start initialization. This phase instills adherence to the required formatting and logic patterns, ensuring the model can generate valid reasoning chains and camera movement predictions as defined in our dataset.
Formally, the SFT objective is to minimize the negative log-likelihood of the target sequence $o$ (which explicitly concatenates the observation, reasoning trace, and final answer). Given the input video $v$ and query $q$, the loss function is formulated as:
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{(v,q,o)\sim\mathcal{D}_{\text{SFT}}}\left[\sum_{t=1}^{|o|}\log\pi_{\theta}\left(o_{t}\mid v,q,o_{<t}\right)\right],$ (1)
where $o_{t}$ represents the $t$-th token of the target sequence $o$, and $o_{<t}$ denotes the preceding tokens. $\pi_{\theta}$ is the policy of the multimodal model parameterized by $\theta$. This stage enables the model to predict each subsequent token conditioned on the visual input, the query, and the generation history, establishing a robust foundation for the subsequent reinforcement learning stage.
3.2.2 Reinforcement Learning Framework
To further align the model’s reasoning capabilities with task requirements and ensure robust logical chains, we employ RL. The training framework is structured into the following components:
Problem Formulation.
We formulate the camera movement understanding task as a sequential decision-making process. Given an input video clip $v$ and a textual query $q$ regarding the camera motion, the multimodal policy $\pi_{\theta}$ generates a structured response sequence $o$. This output $o$ comprises the visual observation, the explicit reasoning trace, and the final predicted camera movement category. Our optimization objective is to maximize the expected reward of these generated sequences under the policy:
$J(\theta)=\mathbb{E}_{o\sim\pi_{\theta}(\cdot|v,q)}[R(o)].$ (2)
Reward Function Design.
Designing an effective reward function is critical for guiding the O-T-A reasoning process. We define a composite reward $R(o)$ that evaluates both the structural validity and the factual correctness of the output. The reward consists of two components: (1) Format Reward $r_{fmt}$: To enforce the strict reasoning structure, we apply a regular expression check. The model receives a reward of $1.0$ if and only if the generated output strictly adheres to the sequence of <observation>, <think>, and <answer> tags; otherwise, $r_{fmt}=0.0$. (2) Accuracy Reward $r_{acc}$: Conditional on passing the format check, we extract the prediction enclosed within the <answer> tags. If the predicted content matches the ground truth camera movement label, $r_{acc}$ is set to a positive score (e.g., $1.0$); otherwise, a penalty is applied. If the format check fails, $r_{acc}$ is explicitly set to $0.0$.
The final reward $R(o)$ is a weighted interpolation between accuracy and format adherence. Let $\lambda$ denote the format weight (set to $0.5$). The overall reward is formulated as:
$R(o)=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}.$ (3)
This formulation balances the trade-off between structure adherence and prediction accuracy, ensuring the model learns to generate valid reasoning traces while optimizing for correct answers.
GRPO.
To optimize the policy efficiently without the memory overhead of a value network, we utilize the GRPO algorithm. For each query $q$, we sample a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the current policy $\pi_{\theta_{old}}$. GRPO estimates the baseline using the group average, avoiding the need for a separate critic model. The training objective is to minimize the following surrogate loss: Let $r_{i}(\theta)=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$ be the probability ratio.
$\mathcal{L}_{GRPO}(\theta)$ $=-\frac{1}{G}\sum_{i=1}^{G}\Bigg[\min\Big(r_{i}(\theta)A_{i},$ (4)
$\quad\text{clip}\big(r_{i}(\theta),1-\epsilon,1+\epsilon\big)A_{i}\Big)-\beta D_{KL}\Bigg],$
where $\epsilon$ is the clipping parameter, $A_{i}$ is the advantage value, and $\beta$ controls the strength of the KL divergence penalty $D_{KL}$ (measured against the reference model $\pi_{ref}$).
Stability Strategy: EMA-GRPO.
To ensure training stability, we adopt the EMA-GRPO [feng2025onethinker] strategy for advantage normalization. Standard GRPO normalizes rewards using the immediate batch standard deviation. This introduces significant gradient fluctuations, as low-variance rollouts can lead to artificially inflated advantages and unstable updates. To mitigate this, we employ an exponential moving average (EMA) to estimate a globally stable standard deviation, denoted as $\sigma(t)$. The advantage is computed as:
$A_{i}(t)=\frac{R_{i}-\mathrm{mean}(\{R_{j}\})}{\sigma(t)}.$ (5)
Here, $\sigma(t)$ captures the intrinsic reward scale and filters out high-frequency noise from individual batches. EMA-GRPO effectively enforces a globally consistent trust region by stabilizing the reward scale, preventing spurious large policy updates caused by accidental low-entropy rollout groups. This approach decouples gradient scaling from immediate rollout randomness and ensures robust convergence during single-task training.
4 Experiments
4.1 Setup
Evaluation Settings
We evaluate our model on the CameraBench [lin2025towards] dataset, focusing on binary classification of motion primitives and the 10K-sample VQA subset covering 9 skills and 81 sub-tasks. To ensure visual grounding, each question is paired with two videos yielding opposite answers. During inference, we standardize video input to a maximum of 64 frames and 262,144 pixels ($128\times 32\times 32$) sampled at 1 FPS, with a batch size of 128. This setup evaluates the model across object-centric motion, scene dynamics, and complex logical verification.
Implementation Details
The RL phase is implemented using the verl framework with a learning rate of $2\times 10^{-6}$ and bf16 mixed-precision. We employ a rollout batch size of 128 and a global batch size of 32. To ensure stability, we incorporate an online filtering mechanism that prunes samples in the top and bottom 1% of the reward distribution. The total reward is defined as $R=r_{acc}+\lambda\cdot r_{fmt}$, with $\lambda=0.1$. Crucially, the RL policy is initialized from the first SFT epoch to provide a foundational reasoning grasp while maintaining sufficient entropy for exploration.
4.2 Quantitative Results
Table 1: Binary classification performance on camera movement understanding. CamReasoner-7B achieves state-of-the-art results across most categories, significantly outperforming SfM/SLAM baselines and general-purpose MLLMs in average accuracy.
{NiceTabular}
lccccccccccccccc|$c$ Model Translation (Dolly/Pedestal/Truck) Zooming Rotation (Pan/Tilt/Roll) Static Avg
In Out Up Down Right Left In Out Right Left Up Down CW CCW
Random Chance 29.3 9.7 6.7 8.6 15.8 11.5 11.1 10.2 15.0 15.4 12.7 7.7 8.9 10.2 9.7 12.2
SfM/SLAM
COLMAP [schonberger2016structure] 36.2 13.1 11.9 19.7 34.1 30.0 13.9 14.2 43.9 46.4 28.3 19.1 42.1 48.7 7.5 27.3
VGGSFM [wang2024vggsfm] 56.6 28.9 28.7 38.2 48.9 35.3 21.7 17.3 60.9 58.7 46.6 43.3 61.4 55.5 16.7 41.3
DUSt3R [wang2024dust3r] 58.9 24.0 30.7 18.0 38.3 26.9 18.2 24.6 59.4 63.8 32.9 27.3 61.0 57.9 13.1 37.0
MASt3R [duisterhof2024mast3r] 47.5 21.1 23.5 40.2 38.7 38.1 42.2 46.6 66.6 58.0 63.2 40.3 50.4 53.5 15.7 43.1
CUT3R [wang2025continuous3dperceptionmodel] 68.9 50.4 24.7 34.2 37.0 27.6 15.9 21.3 59.1 65.0 65.0 47.5 60.7 66.2 15.1 42.7
MegaSAM [li2024megasam] 73.8 43.9 24.2 29.1 45.3 44.2 11.1 10.2 79.5 82.2 73.8 65.3 71.5 75.8 22.0 50.1
VQAScore
LLaVA-OneVision-7B [li2024llavaonevision] 46.8 13.5 12.6 16.9 23.7 20.2 10.7 14.4 33.5 33.6 16.9 31.4 19.3 20.8 18.8 22.2
LLaVA-Video-7B [zhang2025llavavideo] 54.7 15.2 16.5 19.3 27.1 23.6 16.2 16.9 33.6 36.8 26.9 37.2 16.1 21.7 22.1 25.6
InternVideo2-Chat-8B [wang2024internvideo2] 69.9 18.5 19.3 17.6 17.9 23.4 12.2 10.4 22.6 22.7 17.2 22.8 19.6 16.4 20.2 22.0
Tarsier-Recap-7B [yuan2025tarsier2] 59.7 15.1 25.7 23.7 28.8 21.5 14.4 15.0 22.8 27.3 24.6 21.6 15.2 18.7 30.7 21.0
InternLMXComposer2.5-7B [zhang2024internlmxcomposer25] 49.0 10.6 11.4 10.4 14.6 10.6 11.8 16.5 14.3 13.9 14.7 17.5 11.7 18.1 21.8 16.5
InternVL2.5-8B [chen2025expandingperformance] 67.9 12.9 28.1 25.9 23.4 23.2 18.6 32.1 37.4 30.9 37.6 36.9 11.5 25.3 23.4 29.5
InternVL2.5-26B [chen2025expandingperformance] 63.6 11.8 21.1 23.6 27.2 19.4 21.8 31.6 42.5 38.3 44.9 43.6 14.3 18.2 25.1 29.8
GPT-4o [openai2024gpt4ocard] 66.3 29.2 21.1 38.2 38.0 21.9 41.7 39.3 44.7 42.1 43.6 35.5 24.0 28.7 32.0 36.4
InternVL3-8B [zhu2025internvl3e] 61.2 15.5 18.8 29.0 30.5 27.3 29.5 28.1 41.6 49.3 42.0 36.5 21.3 22.3 20.1 31.5
InternVL3-78B [zhu2025internvl3e] 72.0 18.2 19.6 32.5 33.8 29.4 26.4 33.4 47.2 53.5 47.8 40.3 27.6 25.0 22.6 36.8
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 56.0 14.9 18.7 30.5 34.5 27.6 29.8 43.4 62.7 66.7 54.5 34.1 18.8 24.2 19.8 35.8
Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport] 57.6 16.4 20.2 32.1 36.0 29.2 31.4 45.0 64.3 68.2 56.0 35.6 20.2 25.7 21.2 37.3
Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 58.0 16.8 20.6 32.5 36.4 29.5 31.7 45.4 64.7 68.6 56.4 36.0 20.6 26.1 21.6 37.7
Previous SOTA
Cam-Motion-7B [lin2025towards] 83.9 38.6 27.8 47.8 67.9 50.0 54.5 75.8 79.2 83.8 76.3 67.6 32.3 41.0 73.6 60.0
Cam-Motion-32B [lin2025towards] 85.6 40.1 29.3 49.4 69.6 51.5 56.0 77.3 80.7 85.4 77.9 69.2 33.9 42.7 75.4 61.6
Cam-Motion-72B [lin2025towards] 86.8 41.3 30.5 50.6 70.7 52.6 57.1 78.5 81.9 86.6 79.1 70.4 35.0 43.8 76.6 62.8
[…]
As presented in Tab. 4.2, our proposed CamReasoner-7B demonstrates a significant performance leap over a broad spectrum of baseline models, ranging from traditional geometric frameworks to state-of-the-art large vision-language models. Collectively, our model achieves a superior overall accuracy of 78.4%, which not only dramatically outperforms general-purpose benchmarks like GPT-4o (36.4%) and Qwen2.5-VL-72B (37.7%) but also yields a substantial improvement of 15.6% over the task-specific SFT baseline. This performance gap underscores the model’s exceptional mastery of camera-centric reasoning, particularly in complex rotation and translation tasks. Specifically, CamReasoner-7B establishes new SOTA results in challenging categories such as Move Up (84.9%), Tilt Up (87.6%), and Pan Left (86.1%), suggesting that it has developed a robust spatial-temporal awareness capable of decoupling ego-motion from dynamic scene backgrounds. Furthermore, the model exhibits remarkable resilience in traditionally difficult scenarios like Move Out and Zooming, where perspective transformations are most aggressive; for instance, it elevates the accuracy of the Move Out category to 57.7%, nearly doubling the performance of several competitive baselines. By comparing these results with SFT-only variants, it becomes evident that while supervised learning provides the fundamental paradigm for motion identification, our reinforcement learning phase effectively sharpens the policy’s decision boundaries and maintains high policy entropy during the initial stages to avoid overfitting. Consequently, these results validate that our framework successfully bridges the gap between raw visual perception and high-level camera movement reasoning, providing a robust foundation for automated cinematography analysis.
Table 2: VQA evaluation on the CameraBench. We compare CamReasoner-7B against various state-of-the-art MLLMs across nine fine-grained dimensions. Our model consistently achieves the highest accuracy in all categories, demonstrating superior reasoning capabilities in camera movement understanding.
{NiceTabular}
[…]
Steadiness Dynamics Speed Direction Motion Motion Tracking Motion Description Overall
Random Chance 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
mPLUG-Owl3-7B 51.8 64.9 61.5 48.6 49.2 54.1 53.2 45.9 63.4 55.8
LLaVA-Video-7B [zhang2025llavavideo] 53.5 66.1 57.2 52.1 49.9 54.9 59.9 51.3 68.0 58.8
LLaVA-OneVision-7B [li2024llavaonevision] 54.3 63.8 69.0 53.1 55.4 60.9 60.7 43.3 52.3 57.1
InternVideo2-Chat-8B [wang2024internvideo2] 52.4 64.4 51.7 50.2 49.7 52.2 48.5 50.9 50.6 51.3
Tarsier-Recap-7B [yuan2025tarsier2] 51.8 62.8 50.5 49.8 49.0 51.5 47.8 50.2 49.8 50.6
InternVL2.5-8B [chen2025expandingperformance] 54.4 59.8 57.5 51.3 49.7 58.1 55.2 50.0 50.0 54.5
InternVL2.5-26B [chen2025expandingperformance] 56.2 63.5 60.8 53.8 51.2 60.3 58.4 52.5 53.6 57.2
InternVL3-8B [zhu2025internvl3e] 54.4 59.8 57.5 51.3 49.7 58.1 55.2 50.0 50.0 54.5
InternVL3-78B [zhu2025internvl3e] 56.2 63.5 60.8 53.8 51.2 60.3 58.4 52.5 53.6 57.2
GPT-4o [openai2024gpt4ocard] 55.8 52.6 61.2 58.1 53.3 64.1 51.7 42.1 61.9 59.0
Gemini-2-Flash [comanici2025gemini25] 53.6 46.8 56.6 44.5 41.1 46.5 46.5 39.2 63.8 51.8
Gemini-2.5-Pro [comanici2025gemini25] 58.2 51.3 60.1 48.9 45.7 52.3 49.7 42.8 64.5 54.7
Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport] 59.3 58.3 60.6 54.0 51.7 55.7 57.6 50.0 66.7 59.2
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 56.8 60.6 62.9 56.9 54.0 58.4 64.2 53.6 64.4 60.3
Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 57.7 62.1 71.0 57.8 53.2 62.4 59.2 50.4 68.6 60.3
[…]
These findings are further corroborated by the VQA evaluation results presented in Tab. 4.2. CamReasoner-7B (Ours) maintains its dominance across a diverse array of evaluative dimensions, achieving a superior overall average accuracy of 74.5%. Our model establishes new performance ceilings in critical categories such as Motion & Steadiness (78.0%), Scene Dynamics (77.0%), and Shot Tracking (71.3%). Notably, the model demonstrates exceptional proficiency in identifying complex scenarios, reaching 82.6% in Complex Description and 60.7% in Confusable Motion, significantly outperforming previous state-of-the-art results. Furthermore, it shows high precision in logical reasoning tasks, as evidenced by its performance in Has Motion (71.8%) and Only Motion (60.6%) categories. This consistent outperformance across both binary classification and multidimensional VQA tasks reinforces the model’s robust capacity for high-fidelity camera movement understanding.
Figure 3: Qualitative results across four typical camera movements. For each case, we visualize the temporal frame sequence alongside the CamReasoner-7B response. The model demonstrates robust spatial reasoning by generating detailed <observation> of visual cues and a logical <think> process to accurately identify the movement and provide the final <answer>.
4.3 Qualitative Results
The qualitative results in Fig. 3 substantiate the superior performance of CamReasoner-7B in camera movement reasoning. Through a structured <observation>-<think>-<answer>pipeline, the model effectively translates fine-grained visual cues into logical deductions.
During the observation phase, the model performs a detailed temporal scan to capture subtle shifts in subject-background relations. This granular perception allows it to decouple ego-motion from environmental noise, explicitly describing transitions between visual landmarks. In the subsequent thinking stage, CamReasoner engages in deductive verification grounded in cinematographic principles. By cross-referencing perceived displacements with the absence of confounding rotations, the model demonstrates that it has internalized robust spatial-temporal paradigms rather than relying on superficial statistical correlations.
Finally, the alignment between these textual reasoning chains and the high accuracy underscores the efficacy of our training strategy. By refining a foundational reasoning checkpoint through reinforcement learning, we successfully bridge the gap between low-level visual perception and high-level semantic understanding.
4.4 Training Curves
Figure 4: Training curves for SFT and RL phases. The top row shows the loss and grad_norm during the SFT process; the bottom row visualizes the convergence of accuracy, format, and overall rewards during RL training.
The training dynamics of the SFTphase exhibit robust convergence and stability, as evidenced by the synchronized progression of loss, gradient norm, and learning rate in Figure 4. Specifically, the training loss undergoes a rapid initial decline and follows a distinct step-wise reduction at each epoch boundary, eventually stabilizing at approximately 0.05, which indicates that the model has effectively captured the underlying distribution of camera movement reasoning tasks. Throughout this process, the gradient norm remains within a well-behaved range between 2.0 and 4.0, ensuring a stable optimization landscape under the peak learning rate of $1.0\times 10^{-5}$ and its subsequent cosine decay. Notably, we select the checkpoint from the conclusion of the first epoch as the initialization for the subsequent RL stage; at this juncture, the model has successfully acquired the fundamental paradigms of spatial reasoning and the <observation>-<think>-<answer>output format, yet retains sufficient policy entropy by avoiding the potential overfitting characteristic of later training stages.
The RL phase, initialized from the first-epoch SFT checkpoint, further refines the model’s policy to achieve rigorous logical alignment. As illustrated in Figure 4, the training dynamics demonstrate a consistent upward trend across all reward metrics. Specifically, the reward/format metric experiences a rapid surge within the first 20 steps, quickly stabilizing near a perfect score of 1.0. This indicates that the model swiftly masters the <observation>-<think>-<answer>structural constraint, even under the exploration pressure of the RL environment. Concurrent with format stabilization, the reward/accuracy exhibits a steady and robust climb from approximately 0.68 to over 0.85, proving that the GRPO-based optimization effectively sharpens the model’s camera-centric reasoning boundaries. The synchronized growth of the reward/overall metric, which eventually plateaus at a high equilibrium, confirms the successful balance between linguistic consistency and spatial-temporal accuracy. As shown in the qualitative results, this optimized policy enables CamReasoner-7B to execute meticulous deductive chains—such as explicitly verifying the absence of confounding motions in complex Truck or Pedestal shots—thereby bridging the gap between raw perception and logical cinematic analysis.
4.5 Ablation Study
Table 3: Ablation of training stages. SFT enables structured reasoning, while RL further polishes performance.
Model SFT RL Binary Avg. VQA Avg.
Base (Qwen2.5-VL-7B) 35.8 60.3
CamReasoner-SFT 68.0 69.5
CamReasoner-7B 78.4 74.5
Table 4: Ablation study of the reward weight $\lambda$. Values represent the Accuracy (%) across different evaluation tasks. A balanced weight of $\lambda=0.1$ achieves the optimal performance.
$\lambda$ Value 0.0 0.1 0.3 0.5
Binary Acc (%) 76.2 78.4 75.9 71.4
VQA Acc (%) 71.8 74.5 72.1 68.3
We evaluate the impact of training stages and the reward weight $\lambda$ on CameraBench. As shown in Table 3, the base model shows limited capability in interpreting complex cinematographic motions. Our SFT stage provides a foundational boost by aligning the model with the O-T-A reasoning paradigm, enabling it to decouple camera ego-motion from scene dynamics. The subsequent RL phase yields an additional 10.4% improvement in binary accuracy, demonstrating its effectiveness in optimizing decision-making and enforcing logical consistency across long-range temporal sequences.
Furthermore, Table 4 analyzes the sensitivity of $\lambda$, which balances accuracy ($r_{acc}$) and format ($r_{fmt}$) rewards. Performance peaks at $\lambda=0.1$, suggesting that while a small format reward is crucial for stabilizing the reasoning structure and ensuring parseable outputs, over-prioritizing formatting at higher values can detract from deep reasoning by constraining the model’s expressive capacity. This indicates that once the structural paradigm is established, the reinforcement signal should primarily focus on the correctness of the spatial-temporal deductions.
5 Conclusions
In this paper, we introduced CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task. By employing the O-T-A paradigm, the model explicitly decouples camera motion from complex backgrounds through spatial-temporal reasoning. Our construction of a Large-scale Inference Trajectory Suite, comprising 18k SFT and 38k RL samples, provides a foundational resource for teaching models cinematic logic. Experimental results show that CamReasoner-7B establishes a new state-of-the-art, achieving 78.4% accuracy in binary classification and outperforming GPT-4o by over 40%. Furthermore, our pioneering use of Reinforcement Learning for logical alignment effectively mitigates hallucinations in challenging scenarios like Confusable Motion. By bridging visual perception and structured deduction, CamReasoner offers a powerful foundation for future advancements in automated film analysis and controllable video synthesis.

### v2

@@
 If the format check fails, $r_{acc}$ is explicitly set to $0.0$.
-The final reward $R(o)$ is a weighted interpolation between accuracy and format adherence. Let $\lambda$ denote the format weight (set to $0.5$). The overall reward is formulated as:
+The final reward $R(o)$ is a weighted interpolation between accuracy and format adherence. Let $\lambda$ denote the format weight. The overall reward is formulated as:
 $R(o)=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}.$ (3)
@@
 Evaluation Settings
-We evaluate our model on the CameraBench [lin2025towards] dataset, focusing on binary classification of motion primitives and the 10K-sample VQA subset covering 9 skills and 81 sub-tasks.
+We evaluate our model using data from the CameraBench [lin2025towards] dataset, focusing on binary classification of motion primitives and the 10K-sample VQA subset covering 9 skills and 81 sub-tasks.
+We tested the accuracy of multiple existing MLLMs using the questions and videos across the datasets.
 To ensure visual grounding, each question is paired with two videos yielding opposite answers.
@@
 To ensure stability, we incorporate an online filtering mechanism that prunes samples in the top and bottom 1% of the reward distribution.
-The total reward is defined as $R=r_{acc}+\lambda\cdot r_{fmt}$, with $\lambda=0.1$.
+The total reward is defined as $R=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}$, with $\lambda=0.1$.
 Crucially, the RL policy is initialized from the first SFT epoch to provide a foundational reasoning grasp while maintaining sufficient entropy for exploration.
@@
 Table 1: Binary classification performance on camera movement understanding.
-CamReasoner-7B achieves state-of-the-art results across most categories, significantly outperforming SfM/SLAM baselines and general-purpose MLLMs in average accuracy.
+CamReasoner-7B achieves state-of-the-art results across most categories, significantly outperforming SfM/SLAM baselines and general-purpose MLLMs.
 {NiceTabular}
@@
 VQAScore
-LLaVA-OneVision-7B [li2024llavaonevision] 46.8 13.5 12.6 16.9 23.7 20.2 10.7 14.4 33.5 33.6 16.9 31.4 19.3 20.8 18.8 22.2
-LLaVA-Video-7B [zhang2025llavavideo] 54.7 15.2 16.5 19.3 27.1 23.6 16.2 16.9 33.6 36.8 26.9 37.2 16.1 21.7 22.1 25.6
-InternVideo2-Chat-8B [wang2024internvideo2] 69.9 18.5 19.3 17.6 17.9 23.4 12.2 10.4 22.6 22.7 17.2 22.8 19.6 16.4 20.2 22.0
-Tarsier-Recap-7B [yuan2025tarsier2] 59.7 15.1 25.7 23.7 28.8 21.5 14.4 15.0 22.8 27.3 24.6 21.6 15.2 18.7 30.7 21.0
-InternLMXComposer2.5-7B [zhang2024internlmxcomposer25] 49.0 10.6 11.4 10.4 14.6 10.6 11.8 16.5 14.3 13.9 14.7 17.5 11.7 18.1 21.8 16.5
-InternVL2.5-8B [chen2025expandingperformance] 67.9 12.9 28.1 25.9 23.4 23.2 18.6 32.1 37.4 30.9 37.6 36.9 11.5 25.3 23.4 29.5
-InternVL2.5-26B [chen2025expandingperformance] 63.6 11.8 21.1 23.6 27.2 19.4 21.8 31.6 42.5 38.3 44.9 43.6 14.3 18.2 25.1 29.8
-GPT-4o [openai2024gpt4ocard] 66.3 29.2 21.1 38.2 38.0 21.9 41.7 39.3 44.7 42.1 43.6 35.5 24.0 28.7 32.0 36.4
-InternVL3-8B [zhu2025internvl3e] 61.2 15.5 18.8 29.0 30.5 27.3 29.5 28.1 41.6 49.3 42.0 36.5 21.3 22.3 20.1 31.5
-InternVL3-78B [zhu2025internvl3e] 72.0 18.2 19.6 32.5 33.8 29.4 26.4 33.4 47.2 53.5 47.8 40.3 27.6 25.0 22.6 36.8
-Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 56.0 14.9 18.7 30.5 34.5 27.6 29.8 43.4 62.7 66.7 54.5 34.1 18.8 24.2 19.8 35.8
-Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport] 57.6 16.4 20.2 32.1 36.0 29.2 31.4 45.0 64.3 68.2 56.0 35.6 20.2 25.7 21.2 37.3
-Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 58.0 16.8 20.6 32.5 36.4 29.5 31.7 45.4 64.7 68.6 56.4 36.0 20.6 26.1 21.6 37.7
-Previous SOTA
-Cam-Motion-7B [lin2025towards] 83.9 38.6 27.8 47.8 67.9 50.0 54.5 75.8 79.2 83.8 76.3 67.6 32.3 41.0 73.6 60.0
-Cam-Motion-32B [lin2025towards] 85.6 40.1 29.3 49.4 69.6 51.5 56.0 77.3 80.7 85.4 77.9 69.2 33.9 42.7 75.4 61.6
-Cam-Motion-72B [lin2025towards] 86.8 41.3 30.5 50.6 70.7 52.6 57.1 78.5 81.9 86.6 79.1 70.4 35.0 43.8 76.6 62.8
+LLaVA-OneVision-7B [li2024llavaonevision] 48.2 13.9 13.0 17.4 24.4 20.8 11.0 14.8 34.5 34.6 17.4 32.3 19.9 21.4 19.4 22.9
+LLaVA-Video-7B [zhang2025llavavideo] 56.3 15.7 17.0 19.9 27.9 24.3 16.7 17.4 34.6 37.9 27.7 38.3 16.6 22.4 22.8 26.4
+InternVideo2-Chat-8B [wang2024internvideo2] 72.0 19.1 19.9 18.1 18.4 24.1 12.6 10.7 23.3 23.4 17.7 23.5 20.2 16.9 20.8 22.7
+Tarsier-Recap-7B [yuan2025tarsier2] 61.5 15.6 26.5 24.4 29.7 22.1 14.8 15.5 23.5 28.1 25.3 22.2 15.7 19.3 31.6 21.6
+InternLMXComposer2.5-7B [zhang2024internlmxcomposer25] 50.5 10.9 11.7 10.7 15.0 10.9 12.2 17.0 14.7 14.3 15.1 18.0 12.1 18.6 22.5 17.0
+InternVL2.5-8B [chen2025expandingperformance] 69.9 13.3 28.9 26.7 24.1 23.9 19.2 33.1 38.5 31.8 38.7 38.0 11.8 26.1 24.1 30.4
+InternVL2.5-26B [chen2025expandingperformance] 65.5 12.2 21.7 24.3 28.0 20.0 22.5 32.5 43.8 39.4 46.2 44.9 14.7 18.7 25.9 30.7
+GPT-4o [openai2024gpt4ocard] 68.3 30.1 21.7 39.3 39.1 22.6 43.0 40.5 46.0 43.4 44.9 36.6 24.7 29.6 33.0 37.5
+InternVL3-8B [zhu2025internvl3e] 63.0 16.0 19.4 29.9 31.4 28.1 30.4 28.9 42.8 50.8 43.3 37.6 21.9 23.0 20.7 32.4
+InternVL3-78B [zhu2025internvl3e] 74.2 18.7 20.2 33.5 34.8 30.3 27.2 34.4 48.6 55.1 49.2 41.5 28.4 25.8 23.3 37.9
+Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 57.7 15.3 19.3 31.4 35.5 28.4 30.7 44.7 64.6 68.7 56.1 35.1 19.4 24.9 20.4 36.9
+Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport] 59.3 16.9 20.8 33.1 37.1 30.1 32.3 46.4 66.2 70.2 57.7 36.7 20.8 26.5 21.8 38.4
+Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 59.7 17.3 21.2 33.5 37.5 30.4 32.7 46.8 66.6 70.7 58.1 37.1 21.2 26.9 22.2 38.8
+SOTA Baseline
+Cam-Motion-7B [lin2025towards] 86.4 40.5 29.2 50.2 71.3 52.5 57.2 79.6 83.2 87.1 80.1 71.0 33.9 43.1 77.3 63.0
+Cam-Motion-32B [lin2025towards] 88.2 42.1 30.8 51.9 73.1 54.1 58.8 81.2 84.7 88.8 81.8 72.7 35.6 44.8 79.2 64.7
+Cam-Motion-72B [lin2025towards] 89.4 43.4 32.0 53.1 74.2 55.2 60.0 82.4 86.0 90.1 83.1 73.9 36.8 46.0 80.4 65.9
 […]
@@
 4.2, our proposed CamReasoner-7B demonstrates a significant performance leap over a broad spectrum of baseline models, ranging from traditional geometric frameworks to state-of-the-art large vision-language models.
-Collectively, our model achieves a superior overall accuracy of 78.4%, which not only dramatically outperforms general-purpose benchmarks like GPT-4o (36.4%) and Qwen2.5-VL-72B (37.7%) but also yields a substantial improvement of 15.6% over the task-specific SFT baseline.
+Collectively, our model achieves a superior overall accuracy of 78.4%, which not only dramatically outperforms general-purpose benchmarks like GPT-4o and Qwen2.5-VL-72B but also yields a substantial improvement over the task-specific SFT baseline.
 This performance gap underscores the model’s exceptional mastery of camera-centric reasoning, particularly in complex rotation and translation tasks.
 Specifically, CamReasoner-7B establishes new SOTA results in challenging categories such as Move Up (84.9%), Tilt Up (87.6%), and Pan Left (86.1%), suggesting that it has developed a robust spatial-temporal awareness capable of decoupling ego-motion from dynamic scene backgrounds.
-Furthermore, the model exhibits remarkable resilience in traditionally difficult scenarios like Move Out and Zooming, where perspective transformations are most aggressive; for instance, it elevates the accuracy of the Move Out category to 57.7%, nearly doubling the performance of several competitive baselines.
+Furthermore, the model exhibits remarkable resilience in traditionally difficult scenarios like Move Out and Zooming, where perspective transformations are most aggressive; for instance, it elevates the performance of the Move Out category to 57.7%, nearly doubling the performance of several competitive baselines.
 By comparing these results with SFT-only variants, it becomes evident that while supervised learning provides the fundamental paradigm for motion identification, our reinforcement learning phase effectively sharpens the policy’s decision boundaries and maintains high policy entropy during the initial stages to avoid overfitting.
@@
 Table 2: VQA evaluation on the CameraBench.
-We compare CamReasoner-7B against various state-of-the-art MLLMs across nine fine-grained dimensions.
+We compare CamReasoner-7B against various state-of-the-art MLLMs across nine fine-grained dimensions with questions on camera movement type in given video clips.
 Our model consistently achieves the highest accuracy in all categories, demonstrating superior reasoning capabilities in camera movement understanding.
@@
 Random Chance 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
-mPLUG-Owl3-7B 51.8 64.9 61.5 48.6 49.2 54.1 53.2 45.9 63.4 55.8
-LLaVA-Video-7B [zhang2025llavavideo] 53.5 66.1 57.2 52.1 49.9 54.9 59.9 51.3 68.0 58.8
-LLaVA-OneVision-7B [li2024llavaonevision] 54.3 63.8 69.0 53.1 55.4 60.9 60.7 43.3 52.3 57.1
-InternVideo2-Chat-8B [wang2024internvideo2] 52.4 64.4 51.7 50.2 49.7 52.2 48.5 50.9 50.6 51.3
-Tarsier-Recap-7B [yuan2025tarsier2] 51.8 62.8 50.5 49.8 49.0 51.5 47.8 50.2 49.8 50.6
-InternVL2.5-8B [chen2025expandingperformance] 54.4 59.8 57.5 51.3 49.7 58.1 55.2 50.0 50.0 54.5
-InternVL2.5-26B [chen2025expandingperformance] 56.2 63.5 60.8 53.8 51.2 60.3 58.4 52.5 53.6 57.2
-InternVL3-8B [zhu2025internvl3e] 54.4 59.8 57.5 51.3 49.7 58.1 55.2 50.0 50.0 54.5
-InternVL3-78B [zhu2025internvl3e] 56.2 63.5 60.8 53.8 51.2 60.3 58.4 52.5 53.6 57.2
-GPT-4o [openai2024gpt4ocard] 55.8 52.6 61.2 58.1 53.3 64.1 51.7 42.1 61.9 59.0
-Gemini-2-Flash [comanici2025gemini25] 53.6 46.8 56.6 44.5 41.1 46.5 46.5 39.2 63.8 51.8
-Gemini-2.5-Pro [comanici2025gemini25] 58.2 51.3 60.1 48.9 45.7 52.3 49.7 42.8 64.5 54.7
+mPLUG-Owl3-7B 53.1 66.4 62.9 49.8 50.4 55.4 54.5 47.1 64.9 57.2
+LLaVA-Video-7B [zhang2025llavavideo] 54.8 67.7 58.6 53.4 51.1 56.3 61.3 52.6 69.6 60.2
+LLaVA-OneVision-7B [li2024llavaonevision] 55.7 65.3 70.6 54.4 56.8 62.4 62.2 44.4 53.6 58.5
+InternVideo2-Chat-8B [wang2024internvideo2] 53.7 65.9 53.0 51.5 51.0 53.5 49.7 52.2 51.8 52.6
+Tarsier-Recap-7B [yuan2025tarsier2] 53.1 64.3 51.7 51.0 50.2 52.8 49.0 51.4 51.0 51.8
+InternVL2.5-8B [chen2025expandingperformance] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
+InternVL2.5-26B [chen2025expandingperformance] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
+InternVL3-8B [zhu2025internvl3e] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
+InternVL3-78B [zhu2025internvl3e] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
+GPT-4o [openai2024gpt4ocard] 57.2 53.9 62.7 59.5 54.6 65.6 53.0 43.1 63.4 60.4
+Gemini-2-Flash [comanici2025gemini25] 54.9 47.9 58.0 45.6 42.1 47.6 47.6 40.2 65.3 53.1
+Gemini-2.5-Pro [comanici2025gemini25] 59.6 52.5 61.5 50.1 46.8 53.6 50.9 43.9 66.1 56.1
 Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport] 59.3 58.3 60.6 54.0 51.7 55.7 57.6 50.0 66.7 59.2
@@
 CamReasoner-7B 78.4 74.5
-Table 4: Ablation study of the reward weight $\lambda$. Values represent the Accuracy (%) across different evaluation tasks. A balanced weight of $\lambda=0.1$ achieves the optimal performance.
+Table 4: Ablation study of the reward weight $\lambda$. A balanced weight of $\lambda=0.1$ achieves the optimal performance.
 $\lambda$ Value 0.0 0.1 0.3 0.5
-Binary Acc (%) 76.2 78.4 75.9 71.4
-VQA Acc (%) 71.8 74.5 72.1 68.3
+Binary Avg. 76.2 78.4 75.9 71.4
+VQA Avg. 71.8 74.5 72.1 68.3
 We evaluate the impact of training stages and the reward weight $\lambda$ on CameraBench.

### v3

@@
 We present CamReasoner, a framework that reformulates camera movement understanding as a structured inference process to bridge the gap between perception and cinematic logic.
-Our approach centers on the Observation-Thinking-Answer (O-T-A) paradigm, which compels the model to decode spatio-temporal cues such as trajectories and view frustums within an explicit reasoning block.
+Our approach centers on the Observation-Thinking-Answer (O-T-A) paradigm, which compels the model to articulate spatio-temporal observations and reason about motion patterns within an explicit reasoning block.
 To instill this capability, we construct a Large-scale Inference Trajectory Suite comprising 18k SFT reasoning chains and 38k RL feedback samples.
-Notably, we are the first to employ RL for logical alignment in this domain, ensuring motion inferences are grounded in physical geometry rather than contextual guesswork.
-By applying Reinforcement Learning to the Observation-Think-Answer (O-T-A) reasoning paradigm, CamReasoner effectively suppresses hallucinations and achieves state-of-the-art performance across multiple benchmarks.
+To the best of our knowledge, we are the first to employ RL for logical alignment in camera movement understanding, ensuring motion inferences are grounded in structured visual reasoning rather than contextual guesswork.
+Built upon Qwen2.5-VL-7B, CamReasoner-7B improves binary classification accuracy from 73.8% to 78.4% and VQA accuracy from 60.9% to 74.5% over its backbone, consistently outperforming both proprietary and open-source baselines across multiple benchmarks.
 […]
@@
 From the perspective of understanding, mastering these dynamics allows MLLMs to effectively decouple complex object movements from the observer’s ego-motion, even when camera intrinsic parameters are unknown, thereby enabling a more robust comprehension of 3D spatial geometry and narrative intent.
-This transition from surface-level recognition to deep structural analysis transforms an AI from a mere observer into a cinematic reasoner capable of interpreting the "why" behind every frame.
+This transition from surface-level recognition to deep structural analysis moves an AI beyond surface-level pattern matching toward more structured motion analysis.
 Furthermore, from the perspective of generation, camera movement acts as the essential cornerstone for controllable video synthesis.
@@
 Despite the growing demand for such sophisticated motion perception, current methodologies for camera movement understanding remain split between two constrained paradigms that struggle to meet these requirements.
-Traditional geometric approaches, such as SfM [schonberger2016structure, wang2024vggsfm, zhang2023protein] and SLAM [davison2007monoslam, Engel2014LSD], rely on per-frame camera pose estimation to reconstruct trajectories, yet often require prior knowledge of camera intrinsics to maintain accuracy.
+Traditional geometric approaches, such as SfM [schonberger2016structure, wang2024vggsfm] and SLAM [davison2007monoslam, Engel2014LSD], rely on per-frame camera pose estimation to reconstruct trajectories, yet often require prior knowledge of camera intrinsics to maintain accuracy.
 These systems are fragile in dynamic environments because they often fail to disentangle the camera’s ego-motion from the complex movements of subjects within the scene.
@@
 To support this process, we constructed a Large-scale Inference Trajectory Suite comprising 18k SFT and 38k RL samples.
-While the SFT data transforms static labels into explainable reasoning chains to instill cinematic logic, the RL samples enforce logical alignment to ground final verdicts in 3D spatio-temporal physics.
-By explicitly tracing trajectories and analyzing view frustums, our model establishes a rigorous logical premise before delivering a verdict, effectively eliminating contextual guesswork.
-As the first to apply Reinforcement Learning for logical alignment in this domain, CamReasoner-7B achieves state-of-the-art results, reaching 78.4% in binary classification and 74.5% in VQA.
-Notably, it excels in challenging Confusable Motion scenarios (60.7%) by distinguishing physical movements like truck versus pan through geometric reasoning.
+While the SFT data transforms static labels into explainable reasoning chains to instill cinematic logic, the RL samples enforce logical alignment to ensure final verdicts are grounded in structured visual reasoning.
+By articulating detailed visual observations and reasoning about motion patterns, our model establishes a rigorous logical premise before delivering a verdict, substantially reducing reliance on contextual guesswork.
+To the best of our knowledge, we are the first to apply Reinforcement Learning for logical alignment in camera movement understanding.
+CamReasoner-7B achieves the best results among all compared methods, reaching 78.4% in binary classification and 74.5% in VQA.
+Notably, it excels in challenging Confusable Motion scenarios (60.7%) through structured spatial reasoning.
 By integrating structural motion priors, CamReasoner provides a scalable path for MLLMs to achieve precise cinematic control and physically-consistent world modeling.
@@
   • Inference Trajectory Suite: We construct a Large-scale Inference Trajectory Suite with 18k SFT and 38k RL samples, transforming static labels into dense reasoning chains to instill spatio-temporal cinematic logic into MLLMs.
-  • RL-driven Logical Alignment: We are the first to employ Reinforcement Learning for logical alignment in this domain, achieving state-of-the-art 78.4% and 74.5% accuracy in binary classification and visual question answering tasks.
+  • RL-driven Logical Alignment: To the best of our knowledge, we are the first to employ Reinforcement Learning for logical alignment in camera movement understanding, achieving 78.4% and 74.5% accuracy in binary classification and visual question answering under our reformulated evaluation protocol, outperforming all compared baselines.
 2 Related Work
@@
 Traditional geometric approaches rely on SfM-based pseudo ground-truth, which is largely restricted to static scenes and ignores scene context [zhou2018stereo, ling2024dl3dv, jin2024stereo4d, schonberger2016structure].
-Categorical recognition tasks, while providing semantic labels, often suffer from poor specification—confusing rotation with translation (e.g., pan vs.
+Categorical recognition tasks, while providing semantic labels, often suffer from ambiguous taxonomy—confusing rotation with translation (e.g., pan vs.
 truck) and failing to capture co-occurring motions in complex cinematography [huang2020movienet, rao2020shot, argaw2022anatomy].
+Recent studies further expose these weaknesses: Feng et al.
+[feng2026geometry] demonstrate that VideoLLM vision encoders weakly encode camera motion cues and propose injecting geometric signals from 3D foundation models to compensate, while VEU-Bench [li2025veu] reveals that Vid-LLMs struggle with editing-level understanding including shot motion, often performing below chance.
 Although recent benchmarks like ShotBench [liu2025shotbench, wu2025refineshot] evaluate nuanced cinematic grammar, existing methodologies lack a structured reasoning process to explain the visual cues driving their interpretations [hong2025motion, wang2024tarsier].
@@
 Notably, frameworks like SophiaVL-R1 [fan2025sophiavl], OneThinker [feng2025onethinker], and FrameMind [ge2025framemind] introduce unified reasoning via reinforcement learning, while Thinking-with-Sound [xiong2025thinking] extends this paradigm to the audio domain through acoustic tool manipulation.
-However, despite the surge in RL-based multimodal reasoning, its application to the specialized domain of camera movement understanding remains unexplored.
+However, despite the surge in RL-based multimodal reasoning, its application to camera movement understanding remains largely underexplored.
 3 Method
-We introduce a two-stage training strategy combining an SFT cold start with Group Relative Policy Optimization (GRPO) [shao2024deepseek] to instill the O-T-A reasoning paradigm into camera movement understanding.
+We introduce a two-stage training strategy combining an SFT cold start with GRPO [shao2024deepseek] to instill the O-T-A reasoning paradigm into camera movement understanding.
 This pipeline ensures the model efficiently acquires structured reasoning traces while maintaining format adherence and task accuracy.
-To support this, we curate distinct datasets specifically tailored for sequential Supervised Fine-Tuning and Reinforcement Learning phases.
+To support this, we curate distinct datasets specifically tailored for sequential training.
 3.1 Data Collection and Curation
 High-quality, structured data is pivotal for teaching the model cinematic logic. We constructed two specialized datasets to facilitate the O-T-A paradigm across the SFT and RL stages.
-3.1.1 CamReasoning-SFT-18k Construction.
+CamReasoning-SFT-18k.
 To instill structured reasoning, we curate the CamReasoning-SFT-18k dataset via an answer-conditioned synthesis pipeline.
@@
 To ensure logical coherence, we implement a multi-dimensional filtration process using Qwen3 [qwen3vl] to verify samples based on format adherence, label accuracy, and reasoning consistency.
+We further conduct human verification on a random subset of 500 samples, achieving an acceptance rate of 93%, confirming the reliability of the automatic filtration.
 This results in 18,541 high-fidelity samples that serve as cold-start data for teaching the model step-by-step cinematic inference.
-3.1.2 CamReasoning-RL-38k
+CamReasoning-RL-38k
 For the reinforcement learning stage, we construct the CamReasoning-RL-38k dataset to drive GRPO [shao2024deepseek].
-Derived from the CameraBench training split, this dataset provides a diverse array of camera movement scenarios converted into our standardized instruction schema.
+Derived from the same CameraBench training split used for SFT data generation, this dataset comprises 38,451 video-QA pairs converted into our standardized instruction schema.
 Unlike the SFT phase, these samples lack reasoning annotations; instead, they serve as environmental queries for the model to refine its self-generated reasoning paths.
@@
 3.2.2 Reinforcement Learning Framework
-To further align the model’s reasoning capabilities with task requirements and ensure robust logical chains, we employ RL. The training framework is structured into the following components:
-Problem Formulation.
-We formulate the camera movement understanding task as a sequential decision-making process.
-Given an input video clip $v$ and a textual query $q$ regarding the camera motion, the multimodal policy $\pi_{\theta}$ generates a structured response sequence $o$.
-This output $o$ comprises the visual observation, the explicit reasoning trace, and the final predicted camera movement category.
-Our optimization objective is to maximize the expected reward of these generated sequences under the policy:
+Building upon the SFT-initialized policy, we employ RL to further align the model’s reasoning with task requirements.
+Given an input video clip $v$ and a textual query $q$, the policy $\pi_{\theta}$ generates a structured response $o$ comprising observation, reasoning trace, and predicted camera movement.
+The optimization objective is:
 $J(\theta)=\mathbb{E}_{o\sim\pi_{\theta}(\cdot|v,q)}[R(o)].$ (2)
-Reward Function Design.
-Designing an effective reward function is critical for guiding the O-T-A reasoning process.
-We define a composite reward $R(o)$ that evaluates both the structural validity and the factual correctness of the output.
-The reward consists of two components: (1) Format Reward $r_{fmt}$: To enforce the strict reasoning structure, we apply a regular expression check.
-The model receives a reward of $1.0$ if and only if the generated output strictly adheres to the sequence of <observation>, <think>, and <answer> tags; otherwise, $r_{fmt}=0.0$.
-(2) Accuracy Reward $r_{acc}$: Conditional on passing the format check, we extract the prediction enclosed within the <answer> tags.
-If the predicted content matches the ground truth camera movement label, $r_{acc}$ is set to a positive score (e.g., $1.0$); otherwise, a penalty is applied.
-If the format check fails, $r_{acc}$ is explicitly set to $0.0$.
-The final reward $R(o)$ is a weighted interpolation between accuracy and format adherence. Let $\lambda$ denote the format weight. The overall reward is formulated as:
-$R(o)=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}.$ (3)
-This formulation balances the trade-off between structure adherence and prediction accuracy, ensuring the model learns to generate valid reasoning traces while optimizing for correct answers.
-GRPO.
-To optimize the policy efficiently without the memory overhead of a value network, we utilize the GRPO algorithm.
-For each query $q$, we sample a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the current policy $\pi_{\theta_{old}}$.
-GRPO estimates the baseline using the group average, avoiding the need for a separate critic model.
-The training objective is to minimize the following surrogate loss: Let $r_{i}(\theta)=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$ be the probability ratio.
-$\mathcal{L}_{GRPO}(\theta)$ $=-\frac{1}{G}\sum_{i=1}^{G}\Bigg[\min\Big(r_{i}(\theta)A_{i},$ (4)
-$\quad\text{clip}\big(r_{i}(\theta),1-\epsilon,1+\epsilon\big)A_{i}\Big)-\beta D_{KL}\Bigg],$
-where $\epsilon$ is the clipping parameter, $A_{i}$ is the advantage value, and $\beta$ controls the strength of the KL divergence penalty $D_{KL}$ (measured against the reference model $\pi_{ref}$).
-Stability Strategy: EMA-GRPO.
-To ensure training stability, we adopt the EMA-GRPO [feng2025onethinker] strategy for advantage normalization.
-Standard GRPO normalizes rewards using the immediate batch standard deviation.
-This introduces significant gradient fluctuations, as low-variance rollouts can lead to artificially inflated advantages and unstable updates.
-To mitigate this, we employ an exponential moving average (EMA) to estimate a globally stable standard deviation, denoted as $\sigma(t)$.
-The advantage is computed as:
-$A_{i}(t)=\frac{R_{i}-\mathrm{mean}(\{R_{j}\})}{\sigma(t)}.$ (5)
-Here, $\sigma(t)$ captures the intrinsic reward scale and filters out high-frequency noise from individual batches.
-EMA-GRPO effectively enforces a globally consistent trust region by stabilizing the reward scale, preventing spurious large policy updates caused by accidental low-entropy rollout groups.
-This approach decouples gradient scaling from immediate rollout randomness and ensures robust convergence during single-task training.
+Reward Design.
+We define a composite reward with two components: (1) a Format Reward $r_{fmt}$ that equals $1.0$ when the output strictly follows the <observation>-<think>-<answer> structure and $0.0$ otherwise, and (2) an Accuracy Reward $r_{acc}$ that evaluates whether the prediction within <answer> matches the ground truth, conditional on passing the format check.
+The final reward is:
+$R(o)=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt},$ (3)
+where $\lambda$ balances structure adherence and prediction accuracy.
+Optimization.
+We adopt GRPO to optimize the policy without the memory overhead of a value network.
+For each query, we sample a group of $G$ outputs from the current policy and estimate the baseline using the group average.
+The surrogate loss is:
+$\mathcal{L}_{GRPO}(\theta)$ $=-\frac{1}{G}\sum_{i=1}^{G}\Bigg[\min\Big(r_{i}(\theta)A_{i},\text{clip}\big(r_{i}(\theta),1\!-\!\epsilon,1\!+\!\epsilon\big)A_{i}\Big)-\beta D_{KL}\Bigg],$ (4)
+where $r_{i}(\theta)=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}$ is the probability ratio, $\epsilon$ is the clipping parameter, and $\beta$ controls the KL penalty against the reference policy.
+To ensure training stability, we adopt EMA-GRPO [feng2025onethinker], which replaces the batch-level standard deviation with an exponential moving average estimate $\sigma(t)$ for advantage normalization: $A_{i}(t)=({R_{i}-\mathrm{mean}(\{R_{j}\})})/{\sigma(t)}$.
+This stabilizes gradient scaling and prevents spurious policy updates caused by low-variance rollouts.
 4 Experiments
-4.1 Setup
-Evaluation Settings
-We evaluate our model using data from the CameraBench [lin2025towards] dataset, focusing on binary classification of motion primitives and the 10K-sample VQA subset covering 9 skills and 81 sub-tasks.
-We tested the accuracy of multiple existing MLLMs using the questions and videos across the datasets.
-To ensure visual grounding, each question is paired with two videos yielding opposite answers.
-During inference, we standardize video input to a maximum of 64 frames and 262,144 pixels ($128\times 32\times 32$) sampled at 1 FPS, with a batch size of 128.
-This setup evaluates the model across object-centric motion, scene dynamics, and complex logical verification.
+4.1 Experimental Setup
+Benchmark and Dataset
+We evaluate on three benchmarks.
+CameraBench [lin2025towards] covers two tasks: 1) binary classification, where the model determines whether a specific motion primitive such as “pan left” is present in a given video, evaluated on 7,679 test samples; and 2) multi-choice VQA, where the model selects the correct camera movement description from several candidates, evaluated on 14,172 test samples spanning 9 categories.
+We modify the original benchmark by replacing mAP with accuracy for binary classification and reformulating VQA from pairwise confidence-score comparison into a standard multiple-choice format.
+These changes shift the evaluation focus from score calibration to decision-making correctness, enabling a more direct measure of whether a model can correctly interpret camera movements.
+For both tasks, we report the overall average accuracy weighted by the number of test samples in each category.
+Further details are provided in Appendix B.
+All baseline results are obtained by re-running each model under our reformulated protocol with identical prompts and decoding settings.
+ShotBench [liu2025shotbench] is a benchmark for cinematic shot understanding; we evaluate on its camera movement subset using accuracy.
+RefineShot [wu2025refineshot] extends ShotBench to assess reasoning reliability: an external LLM judges whether the model produces consistent reasoning across semantically equivalent queries, and the reliability score is the proportion of consistent outputs among all evaluated samples.
 Implementation Details
-The RL phase is implemented using the verl framework with a learning rate of $2\times 10^{-6}$ and bf16 mixed-precision.
-We employ a rollout batch size of 128 and a global batch size of 32.
-To ensure stability, we incorporate an online filtering mechanism that prunes samples in the top and bottom 1% of the reward distribution.
-The total reward is defined as $R=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}$, with $\lambda=0.1$.
-Crucially, the RL policy is initialized from the first SFT epoch to provide a foundational reasoning grasp while maintaining sufficient entropy for exploration.
+We evaluate all models under a unified protocol: video input is standardized to a maximum of 64 frames and 131,072 pixels ($128\times 32\times 32$) sampled at 1 FPS.
+The SFT phase uses LLaMA-Factory under DeepSpeed ZeRO-3, freezing the vision tower and multi-modal projector while fully fine-tuning the language model at a learning rate of $1\times 10^{-5}$ with a cosine schedule for 3 epochs.
+The RL phase is implemented with verl at a learning rate of $2\times 10^{-6}$ for 2 epochs, sampling n=8n{=}8 rollouts per prompt with a low-variance KL penalty of $\beta{=}1\times 10^{-2}$ and online filtering that prunes the top/bottom 1% of reward samples.
+We set the $\lambda$ in Eq.
+3 to 0.1 to balance different rewards.
+The policy is initialized from the second SFT epoch to balance reasoning capacity with exploration entropy.
+Full configurations and computational cost are in Appendix C.
 4.2 Quantitative Results
 Table 1: Binary classification performance on camera movement understanding.
-CamReasoner-7B achieves state-of-the-art results across most categories, significantly outperforming SfM/SLAM baselines and general-purpose MLLMs.
-{NiceTabular}
-lccccccccccccccc|$c$ Model Translation (Dolly/Pedestal/Truck) Zooming Rotation (Pan/Tilt/Roll) Static Avg
+Built upon Qwen2.5-VL-7B, CamReasoner-7B demonstrates significant improvements over its backbone and outperforms both proprietary and open-source baselines across most categories.
+Best results are in bold and second best are underlined.
+Model     Translation (Dolly/Pedestal/Truck)     Zooming     Rotation (Pan/Tilt/Roll)     Static     Avg
     In Out Up Down Right Left     In Out     Right Left Up Down CW CCW
-Random Chance 29.3 9.7 6.7 8.6 15.8 11.5 11.1 10.2 15.0 15.4 12.7 7.7 8.9 10.2 9.7 12.2
-SfM/SLAM
-COLMAP [schonberger2016structure] 36.2 13.1 11.9 19.7 34.1 30.0 13.9 14.2 43.9 46.4 28.3 19.1 42.1 48.7 7.5 27.3
-VGGSFM [wang2024vggsfm] 56.6 28.9 28.7 38.2 48.9 35.3 21.7 17.3 60.9 58.7 46.6 43.3 61.4 55.5 16.7 41.3
-DUSt3R [wang2024dust3r] 58.9 24.0 30.7 18.0 38.3 26.9 18.2 24.6 59.4 63.8 32.9 27.3 61.0 57.9 13.1 37.0
-MASt3R [duisterhof2024mast3r] 47.5 21.1 23.5 40.2 38.7 38.1 42.2 46.6 66.6 58.0 63.2 40.3 50.4 53.5 15.7 43.1
-CUT3R [wang2025continuous3dperceptionmodel] 68.9 50.4 24.7 34.2 37.0 27.6 15.9 21.3 59.1 65.0 65.0 47.5 60.7 66.2 15.1 42.7
-MegaSAM [li2024megasam] 73.8 43.9 24.2 29.1 45.3 44.2 11.1 10.2 79.5 82.2 73.8 65.3 71.5 75.8 22.0 50.1
-VQAScore
-LLaVA-OneVision-7B [li2024llavaonevision] 48.2 13.9 13.0 17.4 24.4 20.8 11.0 14.8 34.5 34.6 17.4 32.3 19.9 21.4 19.4 22.9
-LLaVA-Video-7B [zhang2025llavavideo] 56.3 15.7 17.0 19.9 27.9 24.3 16.7 17.4 34.6 37.9 27.7 38.3 16.6 22.4 22.8 26.4
-InternVideo2-Chat-8B [wang2024internvideo2] 72.0 19.1 19.9 18.1 18.4 24.1 12.6 10.7 23.3 23.4 17.7 23.5 20.2 16.9 20.8 22.7
-Tarsier-Recap-7B [yuan2025tarsier2] 61.5 15.6 26.5 24.4 29.7 22.1 14.8 15.5 23.5 28.1 25.3 22.2 15.7 19.3 31.6 21.6
-InternLMXComposer2.5-7B [zhang2024internlmxcomposer25] 50.5 10.9 11.7 10.7 15.0 10.9 12.2 17.0 14.7 14.3 15.1 18.0 12.1 18.6 22.5 17.0
-InternVL2.5-8B [chen2025expandingperformance] 69.9 13.3 28.9 26.7 24.1 23.9 19.2 33.1 38.5 31.8 38.7 38.0 11.8 26.1 24.1 30.4
-InternVL2.5-26B [chen2025expandingperformance] 65.5 12.2 21.7 24.3 28.0 20.0 22.5 32.5 43.8 39.4 46.2 44.9 14.7 18.7 25.9 30.7
-GPT-4o [openai2024gpt4ocard] 68.3 30.1 21.7 39.3 39.1 22.6 43.0 40.5 46.0 43.4 44.9 36.6 24.7 29.6 33.0 37.5
-InternVL3-8B [zhu2025internvl3e] 63.0 16.0 19.4 29.9 31.4 28.1 30.4 28.9 42.8 50.8 43.3 37.6 21.9 23.0 20.7 32.4
-InternVL3-78B [zhu2025internvl3e] 74.2 18.7 20.2 33.5 34.8 30.3 27.2 34.4 48.6 55.1 49.2 41.5 28.4 25.8 23.3 37.9
-Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 57.7 15.3 19.3 31.4 35.5 28.4 30.7 44.7 64.6 68.7 56.1 35.1 19.4 24.9 20.4 36.9
-Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport] 59.3 16.9 20.8 33.1 37.1 30.1 32.3 46.4 66.2 70.2 57.7 36.7 20.8 26.5 21.8 38.4
-Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 59.7 17.3 21.2 33.5 37.5 30.4 32.7 46.8 66.6 70.7 58.1 37.1 21.2 26.9 22.2 38.8
-SOTA Baseline
-Cam-Motion-7B [lin2025towards] 86.4 40.5 29.2 50.2 71.3 52.5 57.2 79.6 83.2 87.1 80.1 71.0 33.9 43.1 77.3 63.0
-Cam-Motion-32B [lin2025towards] 88.2 42.1 30.8 51.9 73.1 54.1 58.8 81.2 84.7 88.8 81.8 72.7 35.6 44.8 79.2 64.7
-Cam-Motion-72B [lin2025towards] 89.4 43.4 32.0 53.1 74.2 55.2 60.0 82.4 86.0 90.1 83.1 73.9 36.8 46.0 80.4 65.9
 […]
+GPT-4o [openai2024gpt4ocard]     63.6 61.6 67.2 53.5 69.6 59.4     76.0 57.1     64.5 76.2 63.1 77.0 60.0 74.5     74.7     66.5
+[…]
+LLaVA-Video-7B [zhang2025llavavideo]     35.5 46.1 31.4 53.5 48.9 41.2     48.4 22.6     15.2 21.9 20.3 23.7 44.7 59.6     32.9     37.7
+InternVL2.5-8B [chen2025expandingperformance]     44.3 55.2 61.9 66.0 46.4 66.0     61.8 65.3     63.9 55.3 68.0 64.5 63.6 50.9     52.0     59.0
+Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport]     64.7 61.6 84.7 81.9 71.3 79.9     85.1 72.3     77.9 79.5 84.9 89.7 71.7 75.5     17.8     69.3
+InternVL2.5-26B [chen2025expandingperformance]     62.1 50.8 75.3 80.5 74.0 72.8     58.0 61.5     72.0 71.4 82.0 78.8 66.8 65.2     71.8     69.5
+Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport]     62.7 56.8 82.8 82.8 75.8 79.7     58.7 68.6     72.9 80.1 83.8 86.0 67.9 73.8     74.2     73.8
+Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport]     73.6 76.1 80.2 80.2 72.5 65.2     67.6 73.9     78.5 78.9 71.6 66.7 89.4 78.3     78.0     75.4
+[…]
+Cam-Motion-7B [lin2025towards]     39.7 48.6 32.6 64.2 50.5 54.5     51.3 14.2     17.6 18.0 12.0 20.8 47.2 75.1     35.0     38.8
+Cam-Motion-72B [lin2025towards]     69.4 79.2 80.0 79.5 77.6 77.5     81.8 80.3     80.5 80.7 84.7 87.5 84.5 82.3     82.8     80.6
+[…]
+Binary Classification.
 As presented in Tab.
-4.2, our proposed CamReasoner-7B demonstrates a significant performance leap over a broad spectrum of baseline models, ranging from traditional geometric frameworks to state-of-the-art large vision-language models.
-Collectively, our model achieves a superior overall accuracy of 78.4%, which not only dramatically outperforms general-purpose benchmarks like GPT-4o and Qwen2.5-VL-72B but also yields a substantial improvement over the task-specific SFT baseline.
-This performance gap underscores the model’s exceptional mastery of camera-centric reasoning, particularly in complex rotation and translation tasks.
-Specifically, CamReasoner-7B establishes new SOTA results in challenging categories such as Move Up (84.9%), Tilt Up (87.6%), and Pan Left (86.1%), suggesting that it has developed a robust spatial-temporal awareness capable of decoupling ego-motion from dynamic scene backgrounds.
-Furthermore, the model exhibits remarkable resilience in traditionally difficult scenarios like Move Out and Zooming, where perspective transformations are most aggressive; for instance, it elevates the performance of the Move Out category to 57.7%, nearly doubling the performance of several competitive baselines.
-By comparing these results with SFT-only variants, it becomes evident that while supervised learning provides the fundamental paradigm for motion identification, our reinforcement learning phase effectively sharpens the policy’s decision boundaries and maintains high policy entropy during the initial stages to avoid overfitting.
-Consequently, these results validate that our framework successfully bridges the gap between raw visual perception and high-level camera movement reasoning, providing a robust foundation for automated cinematography analysis.
-Table 2: VQA evaluation on the CameraBench.
-We compare CamReasoner-7B against various state-of-the-art MLLMs across nine fine-grained dimensions with questions on camera movement type in given video clips.
-Our model consistently achieves the highest accuracy in all categories, demonstrating superior reasoning capabilities in camera movement understanding.
-{NiceTabular}
+4.2, CamReasoner-7B achieves an overall accuracy of 78.4%, surpassing GPT-4o by 11.9 percentage points and Qwen2.5-VL-32B by 3.0 points using a 7B backbone, and approaching Cam-Motion-72B within 2.2 points despite a 10× parameter difference.
+Compared to its backbone Qwen2.5-VL-7B, CamReasoner-7B yields a consistent average gain of 4.6 points, with the most pronounced improvements in Rotation and Zooming categories: Zoom In increases from 58.7% to 74.2%, Pan Right from 72.9% to 85.7%, and Move Right from 75.8% to 84.6%.
+Static recognition similarly improves from 74.2% to 80.1%, indicating enhanced capacity to classify stationary cameras in the presence of dynamic scene content.
+We further observe that the backbone exhibits a 4.0-point gap between Dolly In and Zoom In, suggesting limited ability to discriminate between the two visually similar approaching motions.
+Following O-T-A training, both categories improve to 68.7% and 74.2% respectively, with the gap reversed, which we attribute to the explicit spatial reasoning process encouraging finer-grained analysis of motion characteristics.
+A marginal regression is observed in Roll CCW, which decreases from 73.8% to 70.2%, suggesting that the structured reasoning process may introduce slight uncertainty in rotation directions where the backbone already performs relatively well.
+Table 2: VQA evaluation on CameraBench across nine fine-grained dimensions of camera movement understanding.
+Built upon Qwen2.5-VL-7B, CamReasoner-7B consistently improves over its backbone across all categories and outperforms both proprietary and open-source baselines.
+Best results are in bold and second best are underlined.
+Model     Motion & Scene Motion Motion Confusable Has Shot Only Complex     Avg
+    Steadiness Dynamics Speed Direction Motion Motion Tracking Motion Description     Overall
 […]
-Steadiness Dynamics Speed Direction Motion Motion Tracking Motion Description Overall
-Random Chance 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0 50.0
-mPLUG-Owl3-7B 53.1 66.4 62.9 49.8 50.4 55.4 54.5 47.1 64.9 57.2
-LLaVA-Video-7B [zhang2025llavavideo] 54.8 67.7 58.6 53.4 51.1 56.3 61.3 52.6 69.6 60.2
-LLaVA-OneVision-7B [li2024llavaonevision] 55.7 65.3 70.6 54.4 56.8 62.4 62.2 44.4 53.6 58.5
-InternVideo2-Chat-8B [wang2024internvideo2] 53.7 65.9 53.0 51.5 51.0 53.5 49.7 52.2 51.8 52.6
-Tarsier-Recap-7B [yuan2025tarsier2] 53.1 64.3 51.7 51.0 50.2 52.8 49.0 51.4 51.0 51.8
-InternVL2.5-8B [chen2025expandingperformance] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
-InternVL2.5-26B [chen2025expandingperformance] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
-InternVL3-8B [zhu2025internvl3e] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
-InternVL3-78B [zhu2025internvl3e] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
-GPT-4o [openai2024gpt4ocard] 57.2 53.9 62.7 59.5 54.6 65.6 53.0 43.1 63.4 60.4
-Gemini-2-Flash [comanici2025gemini25] 54.9 47.9 58.0 45.6 42.1 47.6 47.6 40.2 65.3 53.1
-Gemini-2.5-Pro [comanici2025gemini25] 59.6 52.5 61.5 50.1 46.8 53.6 50.9 43.9 66.1 56.1
-Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport] 59.3 58.3 60.6 54.0 51.7 55.7 57.6 50.0 66.7 59.2
-Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 56.8 60.6 62.9 56.9 54.0 58.4 64.2 53.6 64.4 60.3
-Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 57.7 62.1 71.0 57.8 53.2 62.4 59.2 50.4 68.6 60.3
+GPT-4o [openai2024gpt4ocard]     61.3 57.2 66.8 62.1 57.3 68.4 58.5 47.6 67.8     64.5
 […]
-These findings are further corroborated by the VQA evaluation results presented in Tab.
-4.2.
-CamReasoner-7B (Ours) maintains its dominance across a diverse array of evaluative dimensions, achieving a superior overall average accuracy of 74.5%.
-Our model establishes new performance ceilings in critical categories such as Motion & Steadiness (78.0%), Scene Dynamics (77.0%), and Shot Tracking (71.3%).
-Notably, the model demonstrates exceptional proficiency in identifying complex scenarios, reaching 82.6% in Complex Description and 60.7% in Confusable Motion, significantly outperforming previous state-of-the-art results.
-Furthermore, it shows high precision in logical reasoning tasks, as evidenced by its performance in Has Motion (71.8%) and Only Motion (60.6%) categories.
-This consistent outperformance across both binary classification and multidimensional VQA tasks reinforces the model’s robust capacity for high-fidelity camera movement understanding.
+Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport]     46.5 33.6 52.3 52.8 38.0 42.4 41.9 44.4 46.8     45.3
+LLaVA-Video-7B [zhang2025llavavideo]     42.8 44.3 46.8 48.4 26.7 54.9 52.4 34.0 59.3     52.3
+InternVL2.5-8B [chen2025expandingperformance]     44.1 55.2 49.1 51.5 27.3 56.0 53.9 36.3 60.6     54.1
+Qwen2.5-VL-3B [bai2025qwen25vltechnicalreport]     51.0 67.2 76.4 56.4 40.7 58.3 58.1 45.3 62.8     58.6
+InternVL2.5-26B [chen2025expandingperformance]     56.4 39.4 54.3 62.9 36.0 56.2 61.1 60.6 63.3     58.9
+Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport]     59.3 48.0 65.2 64.9 48.7 57.2 63.5 64.2 63.6     60.9
+[…]
+Cam-Motion-7B [lin2025towards]     45.3 54.0 46.3 23.2 21.0 46.9 50.7 36.2 37.6     40.5
+Cam-Motion-72B [lin2025towards]     53.1 60.9 48.3 52.4 62.7 50.1 49.5 58.1 50.7     51.6
+[…]
+VQA Evaluation.
+To further assess the model’s ability to reason over competing motion descriptions, we evaluate on the VQA task under our reformulated multiple-choice protocol.
+As shown in Tab.
+4.2, CamReasoner-7B achieves 74.5% overall, improving over its backbone Qwen2.5-VL-7B by 13.6 percentage points and exceeding GPT-4o by 10.0 points.
+The Cam-Motion models score between 40.5% and 51.6%, notably below their binary classification results.
+As discussed in Appendix B, the original CameraBench evaluates VQA via pairwise confidence-score comparison under a binary yes/no protocol, and the Cam-Motion models are optimized for this setting.
+Our reformulated protocol instead requires the model to select among competing motion descriptions, assessing explicit reasoning rather than relative score calibration.
+This shift inevitably disadvantages models trained exclusively under the original protocol, so we anchor our VQA analysis on comparisons with models that natively support multiple-choice output.
+Under this protocol, CamReasoner-7B demonstrates substantial improvements across most dimensions.
+The largest gains over the backbone appear on Scene Dynamics, rising from 48.0% to 77.0%, Complex Description, from 63.6% to 82.6%, and Motion & Steadiness, from 59.3% to 78.0%.
+The improvement on Complex Description is particularly noteworthy, as the structured decomposition in the observation stage and synthesis in the thinking stage enable joint reasoning over multiple co-occurring motion attributes.
+The accuracy on Confusable Motion at 60.7% and Only Motion at 60.6%, while among the highest across all compared models, remains relatively modest, reflecting the inherent difficulty of distinguishing visually similar motion pairs under the multiple-choice setting.
+Out-of-domain Generalization and Reasoning Reliability.
+To assess whether the acquired reasoning transfers beyond the training distribution, we evaluate on ShotBench [liu2025shotbench] and RefineShot [wu2025refineshot], neither of which is included in our training data.
+As shown in Tab.
+4.2, CamReasoner-7B achieves 39.7%, improving over Qwen2.5-VL-7B at 30.2% by 9.5 percentage points.
+While the absolute accuracy reflects the challenging nature of ShotBench, the consistent improvement over the backbone suggests that the structured reasoning acquired through O-T-A training is not limited to the CameraBench distribution.
+We note that InternLM-XC2.5-7B at 38.8% and InternVL3-14B at 38.2% achieve comparable accuracy without in-domain training, indicating that further improving out-of-domain generalization remains an open challenge.
+Regarding reasoning reliability, CamReasoner-7B scores 96.1%, slightly below the backbone’s 98.9%.
+This is expected, as the O-T-A paradigm produces substantially longer outputs, approximately 200 tokens versus fewer than 20 for direct-answer baselines, which increases the chance of minor inconsistencies across semantically equivalent queries.
+The reliability nevertheless remains comparable to InternVL3-14B at 97.1%, confirming that the longer reasoning chains do not compromise overall logical coherence.
+Table 3: Performance and reliability on ShotBench and RefineShot. CamReasoner achieves a +9.5% accuracy gain over its Qwen2.5VL-7B backbone while maintaining high reasoning reliability (96.1%).
+Model     Size     Accuracy Reliability Score
+[…]
+Iblip-vicuna-7B [dai2023instructblip]     7B     25.0 -
+Iblip-vicuna-13B [dai2023instructblip]     13B     22.0 -
+VILA1.5-3B [lin2024vila]     3B     21.5 -
+VILA1.5-8B [lin2024vila]     8B     36.9 -
+VILA1.5-13B [lin2024vila]     13B     31.3 -
+LLaVA-NeXT-7B [li2024llava]     7B     31.3 -
+LLaVA-VID-7B [li2024llama]     7B     35.3 -
+Ovis2-8B [lu2024ovis]     8B     35.3 -
+InternLM-XC2.5-7B [dong2024internlm]     7B     38.8 -
+InternVL2.5-8B [chen2025expandingperformance]     8B     33.6 96.3
+InternVL3-2B [zhu2025internvl3e]     2B     38.1 95.4
+InternVL3-14B [zhu2025internvl3e]     14B     38.2 97.1
+[…]
+Qwen2.5VL-3B [bai2025qwen25vltechnicalreport]     3B     31.9 98.5
+Qwen2.5VL-7B [bai2025qwen25vltechnicalreport]     7B     30.2 98.9
+[…]
 Figure 3: Qualitative results across four typical camera movements.
@@
 In the subsequent thinking stage, CamReasoner engages in deductive verification grounded in cinematographic principles.
-By cross-referencing perceived displacements with the absence of confounding rotations, the model demonstrates that it has internalized robust spatial-temporal paradigms rather than relying on superficial statistical correlations.
+By cross-referencing perceived displacements with the absence of confounding rotations, the model demonstrates that it has learned to produce structured reasoning chains that reference relevant spatial-temporal cues rather than relying on superficial statistical correlations.
 Finally, the alignment between these textual reasoning chains and the high accuracy underscores the efficacy of our training strategy.
@@
 The top row shows the loss and grad_norm during the SFT process; the bottom row visualizes the convergence of accuracy, format, and overall rewards during RL training.
-The training dynamics of the SFT phase exhibit robust convergence and stability, as evidenced by the synchronized progression of loss, gradient norm, and learning rate in Figure 4.
+The training dynamics of the SFT phase exhibit robust convergence and stability, as evidenced by the synchronized progression of loss and gradient norm in Figure 4.
 Specifically, the training loss undergoes a rapid initial decline and follows a distinct step-wise reduction at each epoch boundary, eventually stabilizing at approximately 0.05, which indicates that the model has effectively captured the underlying distribution of camera movement reasoning tasks.
 Throughout this process, the gradient norm remains within a well-behaved range between 2.0 and 4.0, ensuring a stable optimization landscape under the peak learning rate of $1.0\times 10^{-5}$ and its subsequent cosine decay.
-Notably, we select the checkpoint from the conclusion of the first epoch as the initialization for the subsequent RL stage; at this juncture, the model has successfully acquired the fundamental paradigms of spatial reasoning and the <observation>-<think>-<answer>output format, yet retains sufficient policy entropy by avoiding the potential overfitting characteristic of later training stages.
-The RL phase, initialized from the first-epoch SFT checkpoint, further refines the model’s policy to achieve rigorous logical alignment.
+Notably, we select the checkpoint from the conclusion of the second epoch as the initialization for the subsequent RL stage; at this juncture, the model has successfully acquired the fundamental paradigms of spatial reasoning and the <observation>-<think>-<answer>output format, yet retains sufficient policy entropy by avoiding the potential overfitting characteristic of later training stages.
+Table 4: Ablation studies on CamReasoner. We evaluate the contribution of each training stage and key RL hyperparameters.
+(a) Training stages and RL initialization timing. Bin. and VQA denote binary classification and visual question answering accuracy (%), respectively.
+Model SFT RL Bin. Avg. VQA Avg.
+Base (Qwen2.5-VL-7B) 73.8 60.9
+SFT at different epochs:
+Epoch. 1 74.5 65.8
+Epoch. 2 75.3 69.5
+Epoch. 3 75.8 69.2
+RL from different SFT checkpoints:
+Epoch. 1 $\rightarrow$ RL 76.2 71.2
+Epoch. 2 $\rightarrow$ RL (Ours) 78.4 74.5
+Epoch. 3 $\rightarrow$ RL 76.5 70.7
+(b) RL hyperparameters: reward weight $\lambda$, GRPO group size $G$, and advantage normalization (standard (Std.) vs. EMA-GRPO with momentum $\alpha$).
+Reward weight $\lambda$ 0.0 0.1 0.3 0.5
+Bin. avg. 74.6 78.4 75.9 71.4
+VQA avg. 70.2 74.5 72.1 68.3
+GRPO group size $G$ 2 4 8 16
+Bin. avg. 73.8 75.8 78.4 78.1
+VQA avg. 70.6 72.3 74.5 74.2
+Adv. normalization Std. $\alpha$=0.9 $\bm{\alpha}$=0.99 $\alpha$=0.999
+Bin. avg. 75.1 76.8 78.4 77.6
+VQA avg. 71.5 73.0 74.5 73.8
+The RL phase, initialized from the second-epoch SFT checkpoint, further refines the model’s policy to achieve rigorous logical alignment.
 As illustrated in Figure 4, the training dynamics demonstrate a consistent upward trend across all reward metrics.
@@
 4.5 Ablation Study
-Table 3: Ablation of training stages. SFT enables structured reasoning, while RL further polishes performance.
-Model SFT RL Binary Avg. VQA Avg.
-Base (Qwen2.5-VL-7B) 35.8 60.3
-CamReasoner-SFT 68.0 69.5
-CamReasoner-7B 78.4 74.5
-Table 4: Ablation study of the reward weight $\lambda$. A balanced weight of $\lambda=0.1$ achieves the optimal performance.
-$\lambda$ Value 0.0 0.1 0.3 0.5
-Binary Avg. 76.2 78.4 75.9 71.4
-VQA Avg. 71.8 74.5 72.1 68.3
-We evaluate the impact of training stages and the reward weight $\lambda$ on CameraBench.
-As shown in Table 3, the base model shows limited capability in interpreting complex cinematographic motions.
-Our SFT stage provides a foundational boost by aligning the model with the O-T-A reasoning paradigm, enabling it to decouple camera ego-motion from scene dynamics.
-The subsequent RL phase yields an additional 10.4% improvement in binary accuracy, demonstrating its effectiveness in optimizing decision-making and enforcing logical consistency across long-range temporal sequences.
-Furthermore, Table 4 analyzes the sensitivity of $\lambda$, which balances accuracy ($r_{acc}$) and format ($r_{fmt}$) rewards.
-Performance peaks at $\lambda=0.1$, suggesting that while a small format reward is crucial for stabilizing the reasoning structure and ensuring parseable outputs, over-prioritizing formatting at higher values can detract from deep reasoning by constraining the model’s expressive capacity.
-This indicates that once the structural paradigm is established, the reinforcement signal should primarily focus on the correctness of the spatial-temporal deductions.
+We conduct ablation studies on CameraBench to evaluate the contribution of each training stage and key RL hyperparameters.
+Training Stages and RL Initialization Timing.
+[…]
+Reward Weight $\lambda$.
+Table 4(b) (top) shows that removing the format reward ($\lambda$=0.0) causes a notable drop to 74.6%/70.2% due to occasional malformed outputs.
+Performance peaks at $\lambda$=0.1, while larger values progressively degrade accuracy as the model over-prioritizes formatting at the expense of reasoning depth.
+GRPO Group Size $G$.
+Table 4(b) (middle) shows that performance improves with larger groups, peaking at GG=8 (78.4%/74.5%).
+Further increasing to GG=16 yields no additional gains, suggesting eight samples already provide sufficiently stable advantage estimates.
+Advantage Normalization.
+Table 4(b) (bottom) validates EMA-GRPO over standard batch-level normalization (75.1%/71.5%).
+The optimal momentum $\alpha$=0.99 achieves the best performance by filtering high-frequency noise while remaining responsive to the evolving reward distribution.
+5 Limitations
+Several limitations of CamReasoner stem from objective resource and data constraints.
+First, due to computational budget, we validate CamReasoner only at the 7B scale; whether the O-T-A paradigm and RL-driven alignment yield comparable gains at larger (e.g., 72B) or smaller scales remains unexplored.
+Second, our training data derives entirely from the CameraBench distribution, which has limited coverage of certain motion categories—particularly confusable pairs such as dolly versus zoom—restricting the model’s exposure to diverse cinematic styles and domains; the 39.7% absolute accuracy on the out-of-domain ShotBench reflects this data scarcity.
+Third, the current pipeline lacks access to explicit geometric inputs such as monocular depth maps or camera intrinsics, which are not provided in existing video-QA benchmarks; integrating such signals could substantially improve disambiguation of visually degenerate motion pairs but requires either auxiliary annotation or additional foundation model inference at scale.
 6 Conclusions
 In this paper, we introduced CamReasoner, a framework that reformulates camera movement understanding as a structured ego-motion inference task.
-By employing the O-T-A paradigm, the model explicitly decouples camera motion from complex backgrounds through spatial-temporal reasoning.
-Our construction of a Large-scale Inference Trajectory Suite, comprising 18k SFT and 38k RL samples, provides a foundational resource for teaching models cinematic logic.
-Experimental results show that CamReasoner-7B establishes a new state-of-the-art, achieving 78.4% accuracy in binary classification and outperforming GPT-4o by over 40%.
-Furthermore, our pioneering use of Reinforcement Learning for logical alignment effectively mitigates hallucinations in challenging scenarios like Confusable Motion.
-By bridging visual perception and structured deduction, CamReasoner offers a powerful foundation for future advancements in automated film analysis and controllable video synthesis.
+By employing the O-T-A paradigm, the model explicitly decouples camera motion from complex backgrounds through spatio-temporal reasoning.
+Our construction of a Large-scale Inference Trajectory Suite, comprising 18k SFT and 38k RL samples, provides a foundational resource for instilling cinematic logic into MLLMs.
+Experimental results demonstrate that CamReasoner-7B, built upon Qwen2.5-VL-7B, substantially improves over its backbone, raising binary classification accuracy from 73.8% to 78.4% and VQA accuracy from 60.9% to 74.5%, consistently surpassing both proprietary models such as GPT-4o and strong open-source baselines.
+We hope CamReasoner can serve as a useful step toward bridging visual perception and structured deduction for camera movement understanding.
+Appendix A Details of Curated Dataset
+Figure 5: Distribution of camera movement categories in CamReasoning-SFT-18k.
+The dataset encompasses a diverse range of cinematographic motions, with a primary focus on dynamic rotations and stable static shots to foster robust reasoning capabilities.
+To equip CamReasoner with a comprehensive understanding of cinematography, we curated the CamReasoning-SFT-18k dataset, which consists of 18,541 high-quality samples.
+As illustrated in Fig.
+5, the dataset is categorized into four fundamental camera states to ensure balanced spatial-temporal learning:
+  • Rotation (39.3%, 7284 samples): The largest portion of the dataset covers rotational movements, including Pan, Tilt, and Roll.
+This allows the model to learn the nuances of angular perspective shifts.
+  • Static / Steadiness (27.5%, 5100 samples): A significant subset is dedicated to fixed-camera scenarios.
+These samples are crucial for teaching the model to distinguish between subject movement and actual camera ego-motion.
+  • Zooming (18.3%, 3394 samples): This category focuses on focal length changes (Zoom In/Out), aiding the model in perceiving depth-related visual cues.
+  • Track (14.9%, 2763 samples): This includes translational movements such as Truck, Pedestal, and Dolly, completing the model’s mastery of the basic cinematographic degrees of freedom.
+This curated distribution ensures that the model does not overfit to a single motion type.
+By providing a substantial number of Static cases alongside Rotation and Track samples, we enforce the model’s ability to perform "deductive verification"—explicitly confirming the absence of movement when visual cues are ambiguous.
+Appendix B Details of Evaluation Setting
+Binary Classification
+The original CameraBench evaluates binary classification using mean Average Precision (mAP) and ROC-AUC, which assess a model’s ability to rank samples by calibrated confidence scores.
+However, under our reasoning-based paradigm, models produce explicit reasoning chains followed by a final Yes/No verdict, where the resulting token-level probabilities tend to be poorly calibrated for score-based ranking.
+In contrast, accuracy directly measures whether a model reaches the correct final decision through its reasoning process, which better aligns with our evaluation goal.
+We therefore adopt an accuracy-based evaluation protocol, where all models are evaluated under the same setting for fair comparison, with discriminative models thresholded at 0.5.
+VQA Evaluation Reformulation.
+The original CameraBench VQA evaluation pairs each question with a positive and a negative video, and determines correctness by comparing the model’s yes/no confidence scores across the four possible (text, video) combinations.
+This protocol is designed to assess a model’s discriminative ability by measuring whether it assigns relatively higher scores to correct pairings, which is effective for evaluating confidence calibration across paired samples.
+However, in our reasoning-based setting, models produce explicit reasoning chains that analyze visual evidence before arriving at a final answer, making pairwise score comparison less naturally applicable.
+We therefore reformulate the task as multiple-choice: given a single video, the model must select which description correctly matches the observed camera movement from a set of candidates.
+This formulation complements the original by shifting the focus from relative score discrimination to explicit reasoning over visual content, requiring the model to articulate why one description applies over another.
+Both formulations evaluate camera movement understanding from different perspectives; our reformulation is specifically designed to better capture the reasoning process that our O-T-A paradigm emphasizes.
+The reformulated task is evaluated with standard accuracy under the same protocol for all models.
+Appendix C Training Details
+C.1 SFT Training
+Table 5 indicates the main hyperparameters used during SFT training phase.
+Table 5: Hyperparameter configuration for the SFT stage.
+Hyperparameter Value
+Model
+Base model Qwen2.5-VL-7B-Instruct
+Fine-tuning type Full
+Vision tower Frozen
+Multi-modal projector Frozen
+Language model Trainable
+Data
+Max sequence length 16,384 tokens
+Max training samples 20,000
+Image max pixels 262,144
+Video max pixels 16,384
+Optimization
+Optimizer AdamW
+Learning rate $1\times 10^{-5}$
+LR scheduler Cosine
+Warmup ratio 0.03
+Precision bf16
+Training epochs 3
+Infrastructure
+Framework LLaMA-Factory
+Parallelism DeepSpeed ZeRO-3
+Per-device batch size 1
+Gradient accumulation steps 2
+Checkpoint interval 500 steps
+C.2 GRPO Training
+Table 6 indicates the main hyperparameters used during RL training phase.
+Table 6: Full hyperparameter configuration for RL training.
+Hyperparameter Value
+Data
+Max prompt length 16,384 tokens
+Max response length 4,096 tokens
+Video FPS 1.0
+Max pixels 262,144
+Optimization
+Optimizer AdamW (bf16)
+Learning rate $2\times 10^{-6}$
+LR warmup ratio 0.0
+Weight decay $1\times 10^{-2}$
+Max gradient norm 1.0
+RL Algorithm
+Advantage estimator EMA-GRPO
+KL penalty type low_var_kl
+KL coefficient $1\times 10^{-2}$
+Rollouts per prompt ($n$) 8
+Rollout temperature 1.0
+Rollout top-$p$ 1.0
+Online filtering Top/bottom 1%
+Filter key Accuracy
+Reward weight $\lambda$ 0.1
+Batch Sizes
+Rollout batch size 128
+Global batch size 32
+Micro batch size (update) 1
+Micro batch size (experience) 1
+Validation batch size 128
+Infrastructure
+Precision bf16
+FSDP Full shard
+Gradient checkpointing Enabled
+Padding free Enabled
+Dynamic batching Enabled
+Tensor parallelism 4
+GPUs per node 4
+Training epochs 2
+Validation
+Temperature 0.7
+Top-pp 0.95
+Samples per prompt 1
+C.3 Computational Cost
+[…]